A store has hard candy that sells for $4.00 per pound, and soft candy that sells for $6.00 per pound. How much of each should the store use to get 60 pounds of a mix that sells for $5.50 per pound?
Answer by lwsshak3(11628)

A store has hard candy that sells for $4.00 per pound, and soft candy that sells for $6.00 per pound. How much of each should the store use to get 60 pounds of a mix that sells for $5.50 per pound?
***
let x=amount of hard candy to use for the mix
60-x=amount of soft candy to use for the mix
..
4x+6(60-x)=5.5*60
4x+360-6x=330
2x=30
x=15
60-x=45
amount of hard candy to use for the mix=15 lbs
amount of soft candy to use for the mix=45 lbs